Determine the volume generated bounded by the curves y= 1/x, 2x - y = 0, x = 10, and the x-axis. Rotated about the y-axis. *


1
Expert's answer
2021-10-28T07:17:58-0400

Explanation:

According to second theorem of Guldino, volume obtined by a rotation of a section bounded by a function f(x)

and the x-axis between a and b is :"V=\pi\int_a^bf^2(x)dx"


In our case, we have 

"V= \pi \int_1^2e^{(2x)}dx=\pi/2(e^4-e^1)"

so it will be: "Y=1\/10"